OpenStudy (anonymous):

the integral of cosx(1+sin^2x)dx I was trying to solve using parts...but I am not getting to the correct answer. Any help appreciated!

OpenStudy (cruffo):

try u-substitution. distribute cosx, then break the integral up \[\int \cos(x)dx + \int \sin^2(x)\cos(x)dx\] fist integral is basic. In the second, let u = sinx

OpenStudy (anonymous):

Oh yea, that worked perfectly! Thanks for the suggestion

OpenStudy (cruffo):

Oh, sorry... that was suppose to be let u = sin(x) so du = cos(x)dx, and \[\int (1+u^2)du\] But glad you got it!

OpenStudy (anonymous):

Yep, got it. Thanks!